3x+2y=18
3x+8y=-12
Substitution method

Answer by lynnlo(4176)

ANSWER:
X=28/3,,Y=-5
I HAVE SET UP AN EXAMPLE FOR YOU
Y=X+1====A
2X+Y=10==B
(1)ADD PART A INTO PART B
2X+X+1=10
3X+1=10============ONE IS FROM THE ONE AND TEN
3X=9===DIVIDE 3 FROM BOTH SIDES
X=3
2(3)+Y=10
6+Y=10==MINUS 6 FROM BOTH SIDES
Y=4
PLUG IN TO CHECK PROOF:
2(3)+4=10
6+4=10
10=10